# SQL-of-Thought: Multi-agentic Text-to-SQL with Guided Error Correction

### v1

@@ 5.4 Token Cost and other Metrics
LLMs are better and faster at NL2SQL Shi et al. (2025); Li et al. (2024), but the API costs are very expensive. Especially in a multi-agent framework or any multiple-pipeline system, the cost for each sample is a lot. Increasing accuracy comes at the cost of compute and expensive processes. A single run on the Spider benchmark takes about 42.58\$ for Claude Opus 3, and 44.2\$ for GPT models. Switching to cheaper models led to a dip in accuracy, but good-performing models were more expensive. We tried to find a middle ground that is both cost-effective and high-performing. With various combinations of Claude and GPT model inference for different agents, we discovered that tasks such as Schema Linking, Query Plan Generation, and Correction Plan Generation required higher reasoning power and failed when using non-reasoning models. This meant that the Subproblem, SQL, and Correction SQL Agents could use non-reasoning models and still perform at par. This approach enabled a hybrid model of Claude Opus for reasoning-intensive agents and GPT-4o for the other agents to create a good middle ground at approximately a \$30 cost for an entire dataset run, achieving 85% Execution Accuracy (for a 100-sample ablation).
Additionally, we also tested SQL-of-Thought on open-source models, such as LLama-3.1-8B-Instruct (Grattafiori et al., 2024) and Qwen2.5-1.5B (Qwen et al., 2025), and similar variations. We found that not only do these models exhibit high latency, taking up to three times longer for evaluation, but they also underperform significantly, achieving only about 45.3% accuracy on the 100-sample set. These models had trouble generating SQL (led to long strings of repeated hallucination), missing column names across different tables, and were overall not suited for use in an NL2SQL framework. Future work can improve the cost and efficacy of such a framework by fine-tuning SLMs (small language models) on individual agentic tasks. For example, an error database can be designed to fine-tune the Correction Loop agents. Spider Yu et al. (2019) provides detailed fields for each query, including whether it uses clauses such as LIMIT, INTERSECT, GROUPBY, etc. These clauses can be used to create datasets for fine-tuning the Subproblem Agent.
@@ 6 Limitations
Our evaluation is limited to the Spider Yu et al. (2019) benchmark and its variants, which may not fully capture the challenges of real-world databases. The error taxonomy, although effective, has not been exhaustively validated across diverse query structures. The multi-agent design also increases inference costs, relying heavily on closed-source reasoning LLMs. Future work can improve both cost and efficacy by fine-tuning small language models (SLMs) for individual agentic tasks, such as building an error database to fine-tune correction loop agents or utilizing Spider’s clause-level annotations (e.g., LIMIT, INTERSECT, GROUPBY) to fine-tune the Subproblem Agent.

### v2

@@ 5.4 Token Cost and other Metrics
LLMs are better and faster at NL2SQL [16, 7], but the API costs are very expensive. Especially in a multi-agent framework or any multiple-pipeline system, the cost for each sample is a lot. Increasing accuracy comes at the cost of compute and expensive processes. A single run on the Spider benchmark takes about 42.58\$ for Claude Opus 3, and 44.2\$ for GPT models. Switching to cheaper models led to a dip in accuracy, but good-performing models were more expensive. We tried to find a middle ground that is both cost-effective and high-performing. With various combinations of Claude and GPT model inference for different agents, we discovered that tasks such as Schema Linking, Query Plan Generation, and Correction Plan Generation required higher reasoning power and failed when using non-reasoning models. This meant that the Subproblem, SQL, and Correction SQL Agents could use non-reasoning models and still perform at par. This approach enabled a hybrid model of Claude Opus for reasoning-intensive agents and GPT-4o for the other agents to create a good middle ground at approximately a \$30 cost for an entire dataset run, achieving 85% Execution Accuracy (for a 100-sample ablation).
Additionally, we also tested SQL-of-Thought on open-source models, such as LLama-3.1-8B-Instruct [6] and Qwen2.5-1.5B [13], and similar variations. We found that not only do these models exhibit high latency, taking up to three times longer for evaluation, but they also underperform significantly, achieving only about 45.3% accuracy on the 100-sample set. These models had trouble generating SQL (led to long strings of repeated hallucination), missing column names across different tables, and were overall not suited for use in an NL2SQL framework. Future work can improve the cost and efficacy of such a framework by fine-tuning SLMs (small language models) on individual agentic tasks. For example, an error database can be designed to fine-tune the Correction Loop agents. Spider [22] provides detailed fields for each query, including whether it uses clauses such as LIMIT, INTERSECT, GROUPBY, etc. These clauses can be used to create datasets for fine-tuning the Subproblem Agent.
@@ 6 Limitations
Our evaluation is limited to the Spider [22] benchmark and its variants, which may not fully capture the challenges of real-world databases. The error taxonomy, although effective, has not been exhaustively validated across diverse query structures. The multi-agent design also increases inference costs, relying heavily on closed-source reasoning LLMs. Future work can improve both cost and efficacy by fine-tuning small language models (SLMs) for individual agentic tasks, such as building an error database to fine-tune correction loop agents or utilizing Spider’s clause-level annotations (e.g., LIMIT, INTERSECT, GROUPBY) to fine-tune the Subproblem Agent.